In one cycle, a freezer uses 792 J of electrical energy in order to remove 1765 J of heat from its freezer compartment at 10.0 ∘F.

A) What is the coefficient of performance of this freezer?

B) How much heat does it expel into the room during this cycle? (J)
Answer #1

  

A)

Coefficient of performance = Qc/ W = Tc / (Th - Tc )

here Tc = 260.93 K or 10 F and Th = 298K

Coefficient of performance = 260.93 / 37.07 = 7.03 or approximately 7

B)

According to law of thermodynemics

Qh = Qc + W

Qh = 1765+792 = 2557 J